Jim Smith has two options in buying a DVD. He could pay cash of $225 or $30 down plus 15 monthly payments of $19. If he pays cash, how much does he save?

Answer by crabbyoldgamer(13)

If he pays cash he pays $225.


If he pays by the installment plan, he pays $30 + 15*19 = 30 + 285 = $315.


He saves $315 - 225 = $90 by paying cash.